# RelaCtrl: Relevance-Guided Efficient Control for Diffusion Transformers

### v4

@@ Abstract
The Diffusion Transformer plays a pivotal role in advancing text-to-image and text-to-video generation, owing primarily to its inherent scalability. However, existing controlled diffusion transformer methods incur significant parameter and computational overheads and suffer from inefficient resource allocation due to their failure to account for the varying relevance of control information across different transformer layers. To address this, we propose the Relevance-Guided Efficient Controllable Generation framework, RelaCtrl, enabling efficient and resource-optimized integration of control signals into the Diffusion Transformer. First, we evaluate the relevance of each layer in the Diffusion Transformer to the control information by assessing the “ControlNet Relevance Score”—i.e., the impact of skipping each control layer on both the quality of generation and the control effectiveness during inference. Based on the strength of the relevance, we then tailor the positioning, parameter scale, and modeling capacity of the control layers to reduce unnecessary parameters and redundant computations. Additionally, to further improve efficiency, we replace the self-attention and FFN in the commonly used copy block with the carefully designed Two-Dimensional Shuffle Mixer (TDSM), enabling efficient implementation of both the token mixer and channel mixer. Both qualitative and quantitative experimental results demonstrate that our approach achieves superior performance with only 15% of the parameters and computational complexity compared to PixArt-$\delta$. Our project homepage is https://360cvgroup.github.io/RelaCtrl/.
Machine Learning, ICML
1 Introduction
@@
The Diffusion Transformer (DiT) (Peebles & Xie, 2023), with its strong scalability and multi-modal alignment capabilities, has significantly advanced the fields of text-to-image and text-to-video generation (like, PixArt-$\alpha$ (Chen et al., 2023), Flux (BlackForestlabs AI, 2024), Stable Diffusion 3 (Stability AI, 2024), CogVideoX (Yang et al., 2024), Sora (Brooks et al., 2024), HunyuanVideo (Li et al., 2024), and Qihoo-T2X (Wang et al., 2024), etc). By leveraging its robust architecture and scalability, the fidelity of the generated results and consistency with the given textual description are dramatically improved. Recent studies, such as PixArt-$\delta$ (Chen et al., 2024) and OminiControl (Tan et al., 2024), focus on controlled text-to-image generation based on the DiT framework, promoting its application in real-world scenarios such as AI-driven content creation and e-commerce shopping.
@@
Figure 1: Effect of skipping a specific position within the ControlNet block on the quality of the generated image. Higher FID and HDD indicate a more significant impact of the skipped layer on the quality of the final results, reflecting a stronger correlation with the generated image quality.
@@
However, current controlled generation methods for DiT face two main shortcomings. Firstly, a significant number of additional parameters and computations are introduced, increasing the burden on training and inference. For example, PixArt-$\delta$ directly duplicates the first half of the network’s blocks (i.e., 13 blocks), resulting in a 50% increase in both the number of parameters and computational complexity. Similarly, the control token concatenation in OminiControl adds only a limited number of parameters but doubles the number of tokens involved in the attention and linear layers, leading to a nearly 70% increase in overall computational complexity. Secondly, the varying relevance of control information across different layers of the network is often overlooked, resulting in inefficient allocation of computational resources. Our experiments on “ControlNet Relevance Score”, in which we trained a controlled generative model by copying all blocks and removing control blocks from different layers during inference, revealed that different layers in DiT exhibit varying levels of relevance for control information. As shown in Fig. 1, this relevance follows a trend of increasing and then decreasing, with higher relevance observed in the front-center layers and lower relevance in the deeper layers of the network, resulting in only a slight performance loss (for a more detailed explanation, please refer to Sec. 2.1). Existing methods neglect this variation and apply uniform settings to all layers that introduce control information, resulting in inefficient allocation of parameters and computational resources, such as redundant parameters or computations in layers with low relevance.
@@
To address the above issues, we propose the Relevance-Guided Efficient Controllable Generation framework (i.e., RelaCtrl) for diffusion transformer, based on control information relevance analysis. Specifically, to achieve efficient utilization of computational resources, we design relevance-guided allocation and steering strategies. Control blocks are placed at locations with high control information relevance, while locations with weak relevance are left without control blocks. To further reduce the number of parameters and computational complexity introduced by the copy control block operation, we design a lightweight Two-Dimensional Shuffle Mixer (TDSM) to replace the self-attention and FFN layers in the copy block. Self-attention plays the role of token mixer and FFN plays the role of channel mixer. Therefore, to efficiently perform the same functions as the token mixer and channel mixer, TDSM first randomly selects a varying number of feature channel groups, then randomly divides the token groups within each channel group, and finally computes the attention within each token-channel group along the token dimension. Theoretical analysis demonstrates that TDSM can overcome the limitations of local grouping and enable non-local modeling in the channel-token dimensions. In addition, we regulate the number of channel division groups in TDSM based on the correlation. In regions with stronger correlation, we reduce the number of channel groups and expand the feature dimensions involved in attention to enhance its modeling capability. The results from multiple controlled experiments demonstrate that our approach achieves superior performance with only a 45M parameter increase (7.4% of PixArt-$\alpha$) and an additional 46.7 GFLOPs (8.6% of PixArt-$\alpha$).
The main contribution of this paper can be summarized as follows:
@@
  • We investigate in detail the relevance of control information across different layers of the network, finding that the shallower layers are more sensitive to the control signal, while the deeper layers exhibit weaker relevance to the control effect.
@@
  • We propose a Two-Dimensional Shuffle Mixer (TDSM), which efficiently replaces the self-attention and FFN in the original copy block by calculating attention within randomly divided channel and token groups. Theoretical analysis demonstrates that this design overcomes the limitations of local group modeling, ensuring efficient token-mixing and channel-mixing.
@@
  • The experimental results across four different conditional guidance tasks and two text-to-image generation models show that RelaCtrl consistently achieves superior performance while maintaining efficiency, validating the generalization of both the relevance-guided strategy and the proposed TDSM.
2 Methods
2.1 DiT-ControlNet Relevance Prior
@@
In prior studies, the majority of studies on ControlNet have centered on the U-Net architecture. However, the DiT framework (Chen et al., 2023), constructed by stacking a sequence of transformer blocks without an explicit encoder-decoder structure, poses significant challenges for direct adaptation to achieve effective controllability (Chen et al., 2024). To tackle this issue, PixArt-${\delta}$ introduced a method that duplicates the first 13 Transformer blocks from the DiT model, integrates the outputs of these copied blocks with their corresponding frozen blocks, and forwards the combined results to subsequent frozen modules for further processing. While this approach has demonstrated qualitatively favorable results, it brings notable drawbacks. Simply duplicating the first half of the frozen modules results in a considerable increase in model parameters and computational overhead, leading to prohibitively high training and inference costs, particularly for high-resolution image generation. Moreover, our observations indicate that different replicated layers in DiT-ControlNet contribute unequally to the overall generation quality and control fidelity. Blindly copying the initial Transformer blocks can introduce unnecessary computational redundancy without proportional performance gains.
Figure 2: The relevance diagram of different layers in the DiT-ControlNet was calculated based on the FID and HDD ranks. The overall trend shows an initial increase followed by a decrease. The selected placement positions of RelaCtrl in PixArt-$\alpha$ are marked with white numbers.
@@
To systematically evaluate the relevance of individual layers within DiT-ControlNet to generation and controlled quality, we trained a fully controlled PixArt-${\alpha}$ network containing 27 replicated modules. During inference, we systematically skip each control block layer and assess its impact on the final generation. For quantitative assessment, we employed the Fréchet Inception Distance (FID) to measure image generation quality and the Hausdorff Distance (HDD) to evaluate control accuracy. These metrics enabled us to analyze the impact of skipping individual blocks from the control branch on overall performance, providing relevant scores for each control block. Finally, we get the ControlNet Relevance Score $CRS$ based on the combination of these two metrics:
@@
Where $F$ and $H$ represent the rank of initial FID and HDD indicators respectively, $i$ shows the index of the control branch block that is removed, and $\min$ and $\max$ denote the minimum and maximum values within the corresponding rank sequence. If ${{F}_{i}}$ or ${{H}_{i}}$ is higher, it indicates that removing the control block with index $i$ significantly affects the final performance, implying that this module is critically important. Using this approach, we performed single-layer deletions across all replicated blocks in the ControlNet model and derived the regularization metrics and qualitative observations presented in Fig. 1. According to the specified formula 1, the relevance distribution of the ControlNet blocks can be obtained. As illustrated in Fig. 2, the numbers below represent the indices of the control blocks, while the intensity of the shading in each rectangle reflects its relevance to the generation and control performance of the model.
Our findings can be summarized as follows. The most critical modules of DiT-ControlNet are concentrated in the early-middle layers (e.g., blocks 5, 6, and 7). In contrast, removing the last few modules results in only a minimal decline in performance. Overall, the ControlNet Relevance Score exhibits a trend of initially increasing and then decreasing, which contrasts with observations from prior studies of large language models (Gromov et al., 2024; Zhong et al., 2024; Men et al., 2024) or the main branches of the original DiT architecture (Lee et al., 2024). This indicates that simply increasing or decreasing the number of replicated front transformer blocks in DiT-ControlNet does not offer an effective trade-off between performance and computational cost. Such an approach risks removing control modules essential for maintaining optimal performance. Consequently, we propose dynamically guiding the placement and design of control modules within the network by ranking each layer in the DiT model’s control branch according to its relevance. This strategy ensures a more targeted and efficient utilization of network resources.
@@
2.2 Overall Architecture
Fig. 3 depicts the overall pipeline of our proposed method. Based on the ranking of ControlNet Relevance Score derived in Sec. 2.1 and further validated through ablation studies, we identified and selected the 11 most critical control positions—ranked by relevance from high to low—for integrating the control modules, as is shown in Fig. 2. With this approach, we achieve control performance comparable to PixArt-${\delta}$, which utilizes 13 copied modules, while reducing the parameter count by approximately 15%. Although this method effectively decreases the model size and computational overhead, significant redundancy remains in the internal design of the control modules. The transformative power of the Transformer architecture, as emphasized by MetaFormer (Yu et al., 2022), lies in its holistic design, wherein the attention mechanism serves as a token mixer by enabling token-level information mixing, while the remaining components, such as the FFN layers, function as channel mixers to facilitate the integration of channel-wise information. To address the substantial redundancy in the FFN layers within the channel mixer (Pires et al., 2023), we introduce a lightweight module, the Relevance-Guided Lightweight Control Block (RGLC), which unifies token mixing and channel mixing into a single operation. Specifically, we replace the attention and FFN layers in the original PixArt Transformer block with a novel Two-Dimensional Shuffle Mixer (TDSM) design, streamlining the architecture for enhanced efficiency. This method facilitates information interaction and modeling across both the token and channel dimensions, significantly reducing the parameter count and computational demands of the replicated blocks.
2.3 Relevance-Guided Lightweight Control Block
The third part of Fig. 3 illustrates the detailed structure of the RGLC Block. The module takes three inputs: the control condition input $c$, the diffusion timesteps embedding $t$ used to compute the weights for feature normalization, and the input $x$ from the corresponding frozen block. To enhance information interaction between the control branch and the frozen backbone network, $x$ is passed through a zero convolution layer and added to the conditional input $c$, producing ${{c}_{in}}$. The resulting ${{c}_{in}}$ is then processed by the Two-Dimensions Shuffle Mixer (TDSM). Following the processing, the output is passed through a zero convolution layer, resulting in ${{c}_{cond}}$ which is added to the main branch to provide control guidance. This process can be formally expressed as follows:
@@ 2.3 Relevance-Guided Lightweight Control Block
Where ZC refers to the operation of zero convolution. To address the additional computational overhead introduced by the self-attention mechanism, TDSM employs locally grouped self-attention with shuffle characteristics. This design significantly reduces the computational complexity of the network while preserving non-local information interactions within the groups. A detailed analysis of this process is provided in Sec. 2.4.
2.4 Two Dimensions Shuffle Mixer
From the perspective of MetaFormer (Yu et al., 2022), the effectiveness of Transformers can be attributed to two key components: the token mixer, implemented via the self-attention mechanism, and the channel mixer, realized through the feed-forward network (FFN) layer. However, studies have revealed that the FFN is often highly redundant despite consuming a significant proportion of the model parameters (Pires et al., 2023). To alleviate the computational burden of the control branch, we propose grouping tokens for computation while employing specific strategies to enhance interaction and modeling capacity across token groups  (Huang et al., 2021; Cao et al., 2024). Departing from traditional shuffle methods that operate exclusively on the token dimension, we extend the token mixer in the Transformer architecture to model non-local interactions within local windows by introducing a novel approach that jointly operates on both the token and channel dimensions. This dual-dimension strategy enables more efficient and effective modeling. Specifically, We begin by performing random channel selection, followed by random shuffling of the input sequence across the 3D dimension space. Afterward, local self-attention calculations are applied. Although the subsequent attention mechanism is confined to a fixed group, the tokens involved may originate outside this group. This operation effectively disrupts the inherent relationships between tokens and introduces the information flow between channels to some extent, thereby breaking the interaction constraints typically imposed by local attention. To provide theoretical validation, we first present the following definition:
Definition 2.1.
Definition 2.2.
(Random Selection and Shuffle Function). $S:{{c}_{in}}\to c_{rs}^{i},i\in[1,n]$ denote the random selection and shuffle function. This function randomly divides the input ${{c}_{in}}\in{{\mathbb{R}}^{H\times W\times D}}$ into $n$ parts along the channel dimension, and then scrambles the elements within each part in the 3D space. As a result, $c_{rs}^{i}\in{{\mathbb{R}}^{H\times W\times{{d}_{i}}}},i\in[1,n]$, where $\sum\limits_{i=1}^{n}{{{d}_{i}}}=D$.
Definition 2.3.
(Interactive Set and Interactive Distance). Consider $c_{rs}^{i}\in{{\mathbb{R}}^{H\times W\times{{d}_{i}}}}$, where we define the set of token pairs within the same local group after the random shuffle operation as ${{I}_{S}}=\{({{t}_{j}},{{t}_{k}})\}$. This set indicates that, after the random shuffle function $S$, the token indices ${{t}_{j}}$ and ${{t}_{k}}$ are placed within the same local group. At this point, the interactive distance between the two tokens can be defined as ${{d}_{S}}$:
It can be proven that the lower bound of $d({{t}_{j}})$ is $\Omega(\frac{\sqrt{2}}{4}(H+W{{d}_{i}}))$ The detailed derivation steps can be found in the Appendix  B.1.
Theorem 2.4.
The lower bound of $d({{t}_{j}})$ is:
Corollary 2.5.
Let $\bar{d}({{t}_{j}})$ denote the average interactive distance in $c_{rs}^{i}$, and it can be proved that $\bar{d}({{t}_{j}})\approx d({{t}_{j}})$, with the detailed proof provided in the Appendix B.2.
Table 1: Quantitative comparisons of different methods on the COCO validation set (Lin et al., 2014). The best results are highlighted in bold, while the second-best results are marked with underline.
Model Method Controllability Quality Text Consistency Controllability Quality Text Consistency
@@ 2.4 Two Dimensions Shuffle Mixer
PixArt-$\alpha$ PixArt-$\delta$ 96.26 21.38 5.508 0.279 98.91 29.22 5.243 0.275
@@
PixArt-$\alpha$ PixArt-$\delta$ 99.69 35.21 5.723 0.283 0.379 35.50 5.668 0.282
@@
According to Corollary 2.5, the average distance that can be captured by grouped attention in TDSM is $\Omega(\frac{\sqrt{2}}{4}(H+W{{d}_{i}}))$, enabling the modeling of non-local interactions. In summary, the random selection in the channel and shuffle operation we introduced disrupts the positional order of tokens. However, since the token order is not explicitly modeled within the self-attention mechanism for visual tasks, this operation does not compromise the effectiveness of the process. Nonetheless, the arrangement of input tokens may affect the semantic information embedded in the latent code during recovery. To resolve this problem, we propose an inverse recovery operation applied to both the token and channel dimensions following the self-attention computation. This overall method with shuffle and recovery is termed the Two Dimensions Shuffle Mixer (TDSM), leverages the capability of this reversible transformation pairs to ensure information preservation during self-attention calculations, thereby enabling efficient non-local information interaction across both the channel and token dimensions.
3 Experiment
3.1 Experiment Setup
@@
Evaluation Metrics.To comprehensively assess the quality of the generated images, we employed multiple evaluation metrics. The Fréchet Inception Distance (FID) (Heusel et al., 2017) and CLIP-Aesthetics score (Schuhmann et al., 2022)measure the visual fidelity of the generated images, while the CLIP Score (Hessel et al., 2021) evaluates text consistency. The control fidelity is explicitly assessed through the distance between the generated and target images. For images guided by HED and Canny conditions, we utilized the Hausdorff Distance (HDD) (Huttenlocher et al., 1993) for evaluation. For depth conditions, we used MSE-depth, and for segmentation map control, we employed the mean Intersection over Union (mIoU).

### v5

@@ Abstract
The Diffusion Transformer plays a pivotal role in advancing text-to-image and text-to-video generation, owing primarily to its inherent scalability. However, existing controlled diffusion transformer methods incur significant parameter and computational overheads and suffer from inefficient resource allocation due to their failure to account for the varying relevance of control information across different transformer layers. To address this, we propose the Relevance-Guided Efficient Controllable Generation framework, RelaCtrl, enabling efficient and resource-optimized integration of control signals into the Diffusion Transformer. First, we evaluate the relevance of each layer in the Diffusion Transformer to the control information by assessing the ControlNet Relevance Score, which measures the impact of skipping each control layer on both the quality of generation and the control effectiveness during inference. Based on the strength of the relevance, we then tailor the positioning, parameter scale, and modeling capacity of the control layers to reduce unnecessary parameters and redundant computations. Additionally, to further improve efficiency, we replace the self-attention and FFN in the commonly used copy block with the carefully designed Two-Dimensional Shuffle Mixer (TDSM), enabling efficient implementation of both the token mixer and channel mixer. Both qualitative and quantitative experimental results demonstrate that our approach achieves superior performance with only 15% of the parameters and computational complexity compared to PixArt-$\delta$.
footnotetext: * Equal contribution.  Project leader.  Corresponding author.
Introduction
@@
The Diffusion Transformer (DiT) (Peebles and Xie 2023), with its strong scalability and multi-modal alignment capabilities, has significantly advanced the fields of text-to-image and text-to-video generation (like, PixArt-$\alpha$ (Chen et al. 2023), Flux (BlackForestlabs AI 2024), Stable Diffusion 3 (Stability AI 2024), CogVideoX (Yang et al. 2024), Sora (Brooks et al. 2024), HunyuanVideo (Li et al. 2024), and Qihoo-T2X (Wang et al. 2024), etc). By leveraging its robust architecture and scalability, the fidelity of the generated results and consistency with the given textual description are dramatically improved. Recent studies, such as PixArt-$\delta$ (Chen et al. 2024) and OminiControl (Tan et al. 2024), focus on controlled text-to-image generation based on the DiT framework, promoting its application in real-world scenarios such as AI-driven content creation (Ma et al. 2025; He et al. 2025; Wang et al. 2025c) and e-commerce shopping (Lu et al. 2025; Bi et al. 2025; Zhang et al. 2025).
@@
Figure 1: Effect of skipping a specific position within the ControlNet on the quality of the generated image. Higher FID and HDD indicate a more significant impact of the skipped layer on the quality of the final results, reflecting a stronger correlation with the generation.
@@
However, current controlled generation methods for DiT face two main shortcomings. Firstly, a significant number of additional parameters and computations are introduced, increasing the burden on training and inference. For example, PixArt-$\delta$ directly duplicates the first half of the network’s blocks (i.e., 13 blocks), resulting in a 50% increase in both the number of parameters and computational complexity. Similarly, the control token concatenation in OminiControl adds only a limited number of parameters but doubles the number of tokens involved in the attention and linear layers, leading to a nearly 70% increase in overall computational complexity. Secondly, the varying relevance of control information across different layers of the network is often overlooked, resulting in inefficient allocation of computational resources. Our experiments on ”ControlNet Relevance Score”, in which we trained a controlled generative model by copying all blocks and removing control blocks from different layers during inference, revealed that different layers in DiT exhibit varying levels of relevance for control information. As shown in Fig. 1, this relevance follows a trend of increasing and then decreasing, with higher relevance observed in the front-center layers and lower relevance in the deeper layers of the network, resulting in only a slight performance loss (for a more detailed explanation, please refer to Sec . DiT-ControlNet Relevance Prior). Existing methods neglect this variation and apply uniform settings to all layers that introduce control information, resulting in inefficient allocation of parameters and computational resources, such as redundant parameters or computations in layers with low relevance.
@@
To address the above issues, we propose the Relevance-Guided Efficient Controllable Generation framework (i.e., RelaCtrl) for the diffusion transformer, based on control information relevance analysis. Specifically, to achieve efficient utilization of computational resources, we design relevance-guided allocation and steering strategies. Control blocks are placed at locations with high control information relevance, while locations with weak relevance are left without control blocks. To further reduce the number of parameters and computational complexity introduced by the copy control block, we design a lightweight Two-Dimensional Shuffle Mixer (TDSM) to replace the self-attention and FFN layers in the copy block. Self-attention plays the role of token mixer, and FFN plays the role of channel mixer. Therefore, to efficiently perform the same functions as the token mixer and channel mixer, TDSM first randomly selects a varying number of feature channel groups, then randomly divides the token groups within each channel group, and finally computes the attention within each token-channel group along the token dimension. Theoretical analysis demonstrates that TDSM can overcome the limitations of local grouping and enable non-local modeling in the channel-token dimensions. In addition, we regulate the number of channel division groups in TDSM based on the correlation. In regions with stronger correlation, we reduce the number of channel groups and expand the feature dimensions involved in attention to enhance its modeling capability. The results from multiple experiments demonstrate that our approach achieves superior performance with only a 45M parameter increase (7.4% of PixArt-$\alpha$) and an additional 46.7 GFLOPs (8.6% of PixArt-$\alpha$).
Our contribution can be summarized as follows:
@@
  • We investigate in detail the relevance of control information across different layers, finding that the shallower layers are more sensitive to the control signal, while the deeper exhibit weaker relevance to the control effect.
@@
  • We propose a Two-Dimensional Shuffle Mixer, which efficiently replaces the self-attention and FFN in the original copy block by calculating attention within randomly divided channel and token groups. Theoretical analysis demonstrates that this design overcomes the limitations of local group modeling, ensuring efficient token mixing and channel mixing.
@@
  • The experimental results across different conditional guidance tasks show that RelaCtrl consistently achieves superior performance while maintaining efficiency, validating the generalization of both the relevance-guided strategy and the proposed TDSM.
Related Works
Diffusion-Based Models
In recent years, diffusion-based methods have garnered significant success in the field of generation (He et al. 2024; Feng et al. 2024; Wang et al. 2025b; Zhang et al. 2024; Bi et al. 2024), particularly in text-to-image (T2I) generation (Guo et al. 2023; Shuai et al. 2024; Chen et al. 2025; Wang et al. 2025a; Li et al. 2023). These methods utilize text embeddings derived from pre-trained language encoders, such as CLIP (Radford et al. 2021), BERT (Devlin 2018), and T5 (Raffel et al. 2020), to generate images with high fidelity and diversity through an iterative denoising process. Recently, the introduction of the latent diffusion model(Rombach et al. 2022) has marked a significant advancement in this field, enhancing the quality and efficiency of generated content. In pursuit of greater scalability and enhanced generation quality, models like DiT (Peebles and Xie 2023; Chen et al. 2023)integrate large-scale Transformer architectures into the diffusion framework, pushing the boundaries of generative performance. Building on this foundation, Flux (BlackForestlabs AI 2024) synthesizes flow-matching (Lipman et al. 2022) and Transformer-based architecture to achieve state-of-the-art performance.
Controllable Generation with Diffusion Models
Controllable generation has emerged as a prominent area of research in diffusion models (Zhang et al. 2023; Qin et al. 2023; Zhao et al. 2024; Chen et al. 2024; Peng et al. 2024; Che et al. 2024). Currently, there are two main approaches to incorporating controllable conditions into image generation: (1) training a large diffusion model from scratch to enable control under multiple conditions, and (2) fine-tuning a lightweight structure while keeping the original pre-trained model frozen. However, the first approach demands significant computational resources, which limits its accessibility for broader dissemination and personal use. In contrast, recent studies have explored the addition of supplementary network structures to pre-trained diffusion models, allowing for control over the generated outputs without the need to retrain the entire model. ControlNet (Zhang et al. 2023) enables image generation that aligns with control information by reproducing specific layers within the network and connecting them to the original layers using zero convolution. Building on this foundation, Uni-Control (Qin et al. 2023) introduces a Mixture-of-Experts (MoE) framework, which unifies control across multiple spatial conditions. ControlNet-XS (Zavadski et al. 2025) further improves the interaction bandwidth and frequency between the control and main branches within the ControlNet architecture, drawing inspiration from principles of feedback control systems. Nonetheless, these approaches are primarily based on the U-Net structure and may not yield the desired results when directly applied to Diffusion Transformers (DiT) without modification (Chen et al. 2024). PixArt-$\delta$ (Chen et al. 2024) proposed a design methodology specifically tailored for DiT, but directly copying the first half of the network results in a 50% increase in both parameter count and computational complexity, resulting in high computational cost and being inconvenient for community research and practical deployment.
Methods
DiT-ControlNet Relevance Prior
@@ Methods
In prior studies, the majority of studies on ControlNet have centered on the U-Net architecture. However, the DiT framework (Chen et al. 2023), constructed by stacking a sequence of transformer blocks without an explicit encoder-decoder structure, poses significant challenges for direct adaptation to achieve effective controllability (Chen et al. 2024). To tackle this issue, PixArt-$\delta$} introduced a method that duplicates the first 13 Transformer blocks from the DiT model, integrates the outputs of these copied blocks with their corresponding frozen blocks, and forwards the combined results to subsequent frozen modules for further processing. While this approach has demonstrated qualitatively favorable results, it brings notable drawbacks. Simply duplicating the first half of the frozen modules results in a considerable increase in model parameters and computational overhead, leading to prohibitively high training and inference costs, particularly for high-resolution image generation. Moreover, our observations indicate that different replicated layers in DiT-ControlNet contribute unequally to the overall generation quality and control fidelity. Blindly copying the initial Transformer blocks can introduce unnecessary computational redundancy without proportional performance gains.
@@
To systematically evaluate the relevance of individual layers within DiT-ControlNet to generation and controlled quality, we trained a fully controlled PixArt-$\alpha$} network containing 27 replicated modules. We systematically skip each control block layer during inference and assess its impact on the final generation. We employed the Fréchet Inception Distance (FID) to measure image generation quality and the Hausdorff Distance (HDD) to evaluate control accuracy for quantitative assessment. These metrics enabled us to analyze the impact of skipping individual blocks from the control branch on overall performance, providing relevant scores for each control block. Finally, we get the ControlNet Relevance Score $CRS$ based on the combination of these two metrics:
@@
Where $F$ and $H$ represent the rank of initial FID and HDD indicators, respectively, $i$ shows the index of the control branch block that is removed, and $\min$ and $\max$ denote the minimum and maximum values within the corresponding rank sequence. If ${F}_{i}$} or ${H}_{i}$} is higher, it indicates that removing the control block with index $i$ significantly affects the final performance, implying that this module is critically important. Using this approach, we performed single-layer deletions across all replicated blocks in the ControlNet model and derived the regularization metrics and qualitative observations presented in Fig. 1. According to the specified formula 1, the relevance distribution of the ControlNet blocks can be obtained.
Our findings can be summarized as follows. The most critical modules of DiT-ControlNet are concentrated in the early-middle layers (e.g., blocks 5, 6, and 7). In contrast, removing the last few modules results in only a minimal decline in performance. Overall, the ControlNet Relevance Score exhibits a trend of initially increasing and then decreasing, which contrasts with observations from prior studies of large language models (Gromov et al. 2024; Zhong et al. 2024; Men et al. 2024) or the main branches of the original DiT architecture (Lee et al. 2024). This indicates that simply increasing or decreasing the number of replicated front transformer blocks in DiT-ControlNet does not offer an effective trade-off between performance and computational cost. Such an approach risks removing control modules essential for maintaining performance. Consequently, we propose dynamically guiding the placement and design of control modules within the network by ranking each layer in the DiT model’s control branch according to its relevance. This strategy ensures a more targeted and efficient utilization of network resources.
@@
Overall Architecture
Fig. 2 depicts the overall pipeline of our proposed method. Based on the ranking of ControlNet Relevance Score derived in Sec . DiT-ControlNet Relevance Prior and further validated through ablation studies, we identified and selected the 11 most critical control positions—ranked by relevance from high to low—for integrating the control modules. With this approach, we achieve control performance comparable to PixArt-$\delta$}, which utilizes 13 copied modules, while reducing the parameter count by approximately 15%. Although this method effectively decreases the model size and computational overhead, significant redundancy remains in the internal design of the control modules. The transformative power of the Transformer architecture, as emphasized by MetaFormer (Yu et al. 2022), lies in its holistic design, wherein the attention mechanism serves as a token mixer by enabling token-level information mixing, while the remaining components, such as the FFN layers, function as channel mixers to facilitate the integration of channel-wise information. To address the substantial redundancy in the FFN layers within the channel mixer (Pires et al. 2023), we introduce a lightweight module, the Relevance-Guided Lightweight Control Block (RGLC), which unifies token mixing and channel mixing into a single operation. Specifically, we replace the attention and FFN layers in the original PixArt Transformer block with a novel Two-Dimensional Shuffle Mixer (TDSM) design, streamlining the architecture for enhanced efficiency. This method facilitates information interaction and modeling across both token and channel dimensions, significantly reducing the replicated blocks’ parameter and computational demands.
Relevance-Guided Lightweight Control Block
The third part of Fig. 2 illustrates the detailed structure of the RGLC Block. The module takes three inputs: the control condition input $c$, the diffusion timesteps embedding $t$ used to calculate the weights for feature normalization, and the input $x$ from the corresponding frozen block. To enhance information interaction between the control branch and the frozen backbone network, $x$ is passed through a zero convolution layer and added to the conditional input $c$, producing ${c}_{in}$}. The resulting ${c}_{in}$} is then processed by the Two-Dimensional Shuffle Mixer (TDSM). Following the processing, the output is passed through a zero convolution layer, resulting in ${c}_{cond}$}, which is added to the main branch to provide control guidance. This process can be formally expressed as follows:
@@
Where ZC refers to the operation of zero convolution. To address the additional computational overhead introduced by the self-attention mechanism, TDSM employs locally grouped self-attention with shuffle characteristics. This design significantly reduces the computational complexity of the network while preserving non-local information interactions within the groups.
Model Method Controllability Quality TC Controllability Quality TC
@@
$\alpha PixArt$-$\delta$ 96.26 21.38 5.508 0.279 98.91 29.22 5.243 0.275
@@
$\alpha PixArt$-$\delta$ 99.69 35.21 5.723 0.283 0.379 35.50 5.668 0.282
@@
Table 1: Quantitative comparisons of different methods on the COCO validation set (Lin et al. 2014). The best results are highlighted in bold, while the second-best results are underlined. TC: Text Consistency, C-SC: CLIP-Score, C-Ae: CLIP-AE.
Two-Dimensional Shuffle Mixer
From the perspective of MetaFormer (Yu et al. 2022), the effectiveness of Transformers can be attributed to two key components: the token mixer, implemented via the self-attention mechanism, and the channel mixer, realized through the feed-forward network (FFN) layer. However, studies have revealed that the FFN is often highly redundant despite consuming a significant proportion of the model parameters (Pires et al. 2023). To alleviate the computational burden of the control branch, we propose grouping tokens for computation while employing specific strategies to enhance interaction and modeling capacity across token groups  (Huang et al. 2021; Cao et al. 2024). Departing from traditional shuffle methods that operate exclusively on the token dimension, we extend the token mixer in the Transformer architecture to model non-local interactions within local windows by introducing a novel approach that jointly operates on both the token and channel dimensions. This dual-dimension strategy enables more efficient and effective modeling. Specifically, we perform random channel selection, followed by random shuffling of the input sequence across the 3D dimension space. Afterward, local self-attention calculations are applied. Although the subsequent attention mechanism is confined to a fixed group, the tokens involved may originate outside this group. This operation effectively disrupts the inherent relationships between tokens and introduces the information flow between channels to some extent, thereby breaking the interaction constraints typically imposed by local attention. To provide theoretical validation, we first present the following definition:
Definition 1.
Definition 2.
(Random Selection and Shuffle Function). $S:{{c}_{in}}\to c_{rs}^{i},i\in[1,n]$ denote the random selection and shuffle function. This function randomly divides the input ${{c}_{in}}\in{{\mathbb{R}}^{H\times W\times D}}$ into $n$ parts along the channel and then scrambles the elements within each part in the 3D space. As a result, $c_{rs}^{i}\in{{\mathbb{R}}^{H\times W\times{{d}_{i}}}},i\in[1,n]$, where $\sum\limits_{i=1}^{n}{{{d}_{i}}}=D$.
Definition 3.
(Interactive Set and Interactive Distance). Consider $c_{rs}^{i}\in{{\mathbb{R}}^{H\times W\times{{d}_{i}}}}$, where we define the set of token pairs within the same local group after the random shuffle operation as ${{I}_{S}}=\{({{t}_{j}},{{t}_{k}})\}$. This set indicates that, after the random shuffle $S$, the token indices ${t}_{j}$} and ${t}_{k}$} are placed within the same local group. At this point, the interactive distance between the two tokens can be defined as ${d}_{S}$}:
It can be proven that the lower bound of $d({{t}_{j}})$ is $\Omega(\frac{\sqrt{2}}{4}(H+W{{d}_{i}}))$ .
Theorem 4.
The lower bound of $d({{t}_{j}})$ is:
Corollary 5.
Let $\bar{d}({{t}_{j}})$ denote the average interactive distance in $c_{rs}^{i}$, and it can be proved that $\bar{d}({{t}_{j}})\approx d({{t}_{j}})$.
@@
According to Corollary 5, the average distance that can be captured by grouped attention in TDSM is $\Omega(\frac{\sqrt{2}}{4}(H+W{{d}_{i}}))$, enabling the modeling of non-local interactions. In summary, the random selection in the channel and the shuffle operation we introduced disrupt the positional order of tokens. However, since the token order is not explicitly modeled within the self-attention mechanism for visual tasks, this operation does not compromise the effectiveness of the process. Nonetheless, the arrangement of input tokens may affect the semantic information embedded in the latent code during recovery. To resolve this problem, we propose an inverse recovery operation applied to both the token and channel dimensions following the self-attention computation. This overall method with shuffle and recovery is termed the Two-Dimensional Shuffle Mixer (TDSM), leveraging the capability of these reversible transformation pairs to ensure information preservation during self-attention calculations, thereby enabling efficient non-local information interaction across both the channel and token dimensions.
Experiment
Experiment Setup
@@
Evaluation Metrics.To comprehensively assess the quality of the generated images, we employed multiple evaluation metrics. The Fréchet Inception Distance (FID) (Heusel et al. 2017) and CLIP-Aesthetics score (C-Ae) (Schuhmann et al. 2022)measure the visual fidelity of the generated images, while the CLIP Score (C-SC) (Hessel et al. 2021) evaluates text consistency. The control fidelity is explicitly assessed through the distance between the generated and target images. For images guided by HED and Canny conditions, we utilized the Hausdorff Distance (HDD) (Huttenlocher et al. 1993) for evaluation. For depth conditions, we used MSE-depth, and for segmentation map control, we employed the mean Intersection over Union (mIoU).